Carmelita Inc., has the following information available:

Costs from Beginning Inventory Costs from current Period
Direct materials $5,800 $25,300
Conversion costs 6,300 158,600

At the beginning of the period, there were 400 units in process that were 40% complete as to conversion costs and 100% complete as to direct materials costs. During the period, 5,500 units were started and completed. Ending inventory contained 400 units that were 32% complete as to conversion costs and 100% complete as to materials costs. The company uses the FIFO process cost method.

The cost of completing a unit during the current period was

a.$37.58

b.$31.32

c.$27.03

d.$46.97
Answer #1

Unit transferred out = 5500+400 = 5900 Units

Equivalent unit of material = 5900 Units

Equivalent unit of conversion = (400*60%+5500*100%+400*32%) = 5868

Cost per equivalent unit of material = 25300/5900 = 4.29

Cost per equivalent unit of conversion = 158600/5868 = 27.03

Cost of completing unit = 4.29+27.03 = $31.32

So answer is b) $31.32